# RoboDuet: Whole-body Legged Loco-Manipulation with Cross-Embodiment Deployment

### v4

@@ IV-C Training Details
We train 4096 parallel agents using the IsaacGym simulator [12]. We train all algorithms for 50,000 iterations across 3 seeds, with the two-stage training comprising 10,000 iterations for stage 1 and 40,000 for stage 2, employing the asymmetric actor-critic framework used in [28]. All neural networks are designed as the Multilayer Perceptron (MLP) with ELU [29] activations for hidden layers. Unlike the cooperative policy, the unified policy merges feature extractors into one, so we double the number of neurons in the hidden layers. All training is performed on NVIDIA RTX 4090 GPUs.
@@ IV-E Metrics
To quantify the performance of the algorithm, we define several metrics: (1) velocity tracking error, measured by the difference between the target commands and actual states for $v_{x}$, $v_{y}$, and $\omega_{yaw}$; (2) position tracking error, determined by the error for $l$, $p$, $y$, and the Euclidean distance $D$; (3) orientation tracking error, calculated based on the angles $\alpha$, $\beta$, $\gamma$, and the quaternion geodesic distance $\zeta$; (4) survival rate, assessed by randomly applying forces of 10 to 20 newtons to the base for 2 seconds and calculating the proportion of robots that maintain a base height above 0.26 m across all samples; (5) solvability, defined as the ratio of collision-free poses to the total number of samples; and (6) workspace, defined as the area of the convex hull formed by all target commands within the tracking error threshold, ensuring no self-collisions throughout the tracking process. A sample is considered successful if the tracking of the end-effector pose meets the thresholds of $D\leq 0.03$ m for distance and $\theta\leq\pi/36$ for orientation, where $\theta$ is the angle between the two vectors obtained by applying the target orientation and end-effector orientation to the same unit vector. The maximum time allowed to reach the target command is 4 seconds, after which the average error over the following 2 seconds is computed.
@@ V-A Simulation Experiments
V-A1 Ablation
To validate the significance of the two-stage training and the cooperative policy mechanism, which are key components of RoboDuet, we conduct sufficient ablation experiments by comparing with Unified, Two-Stage and Cooperated. The results are shown in Table III, which demonstrate that all configurations meet the requirements for stable standing. Although two-stage training does not significantly improve end-effector pose tracking, it demonstrates a notable improvement in resisting external perturbations. Specifically, Two-Stage improves survival rates by 10.93% over the Unified, attributed to the focus of stage 1 on optimizing locomotion. On the other hand, Cooperated and RoboDuet outperform both the Unified and Two-Stage configurations across almost all metrics, indicating that the cooperative policy mechanism effectively decouples multi-task learning, enabling the two policies to collaborate while maintaining focus. Cooperated achieves the best performance in velocity tracking and position tracking, largely due to its lack of the gait prior constraint from stage 1, allowing it to take a more aggressive approach in achieving higher tracking rewards which results a decrease on survival rates. In comparison, RoboDuet maintains comparable end-effector tracking performance while achieving robust locomotion motion ability and offers a larger operational workspace during motion. In summary, by effectively inte- grating cooperative policy and two-stage training, RoboDuet significantly enhances control performance in both tracking accuracy and gait stability, underscoring the essential role of these components.
Figure 4: Comparison of Solvability and Workspace between Floating Base+IK and RoboDuet.
TABLE III: Ablation Results of Different Methods in Still and Move Modes
@@ V-A1 Ablation
Metrics (scaled by $10^{-2}$) Still Move
Baseline Two-Stage Cooperative RoboDuet Baseline Two-Stage Cooperative RoboDuet
Velocity Tracking $\downarrow$ $v_{x}$ (m/s) 0.47$\pm$0.00 0.48$\pm$0.00 0.34$\pm$0.00 0.32$\pm$0.00 13.58$\pm$0.00 14.00$\pm$0.02 10.12$\pm$0.00 12.12$\pm$0.00
$v_{y}$ (m/s) 0.40$\pm$0.00 0.47$\pm$0.00 0.39$\pm$0.00 0.37$\pm$0.00 20.12$\pm$0.01 18.57$\pm$0.05 15.87$\pm$0.00 16.51$\pm$0.00
$\omega_{z}$ (rad/s) 0.32$\pm$0.00 0.34$\pm$0.00 0.32$\pm$0.00 0.32$\pm$0.00 62.58$\pm$0.00 62.76$\pm$0.00 60.67$\pm$0.00 60.95$\pm$0.00
Position Tracking $\downarrow$ $l$ (m) 4.23$\pm$0.00 3.82$\pm$0.00 2.58$\pm$0.00 2.59$\pm$0.00 4.05$\pm$0.00 3.65$\pm$0.00 2.59$\pm$0.00 2.45$\pm$0.00
$p$ (rad) 26.07$\pm$0.07 23.78$\pm$0.23 19.33$\pm$0.04 20.98$\pm$0.03 22.06$\pm$0.02 21.08$\pm$0.14 18.85$\pm$0.04 19.89$\pm$0.02
$y$ (rad) 19.46$\pm$0.04 17.71$\pm$0.03 11.63$\pm$0.00 10.97$\pm$0.00 18.73$\pm$0.04 15.56$\pm$0.01 11.62$\pm$0.00 10.72$\pm$0.00
$D$ (m) 16.21$\pm$0.01 15.00$\pm$0.04 12.17$\pm$0.01 12.87$\pm$0.01 14.50$\pm$0.00 13.46$\pm$0.02 12.05$\pm$0.01 12.26$\pm$0.00
Orientation Tracking $\downarrow$ $\alpha$ (rad) 43.60$\pm$0.07 46.77$\pm$0.09 42.72$\pm$0.11 43.33$\pm$0.02 44.05$\pm$0.11 46.34$\pm$0.04 42.40$\pm$0.12 43.12$\pm$0.02
$\beta$ (rad) 59.17$\pm$0.02 60.72$\pm$0.04 49.79$\pm$0.06 46.66$\pm$0.04 56.13$\pm$0.03 61.34$\pm$0.23 49.52$\pm$0.05 47.43$\pm$0.05
$\gamma$ (rad) 51.95$\pm$0.03 52.70$\pm$0.06 40.74$\pm$0.03 40.15$\pm$0.02 52.20$\pm$0.04 54.67$\pm$0.07 42.00$\pm$0.03 41.82$\pm$0.00
$\zeta$ (-) 50.87$\pm$0.00 52.17$\pm$0.00 48.10$\pm$0.00 48.07$\pm$0.00 50.62$\pm$0.00 52.25$\pm$0.01 48.16$\pm$0.00 48.10$\pm$0.02
Survival Rate (%) $\uparrow$ 84.46$\pm$0.92 95.39$\pm$0.03 93.80$\pm$0.07 96.60$\pm$0.01 92.83$\pm$1.58 99.50$\pm$0.00 99.37$\pm$0.00 99.91$\pm$0.00
Workspace ($m^{3}$) $\uparrow$ 42.40$\pm$0.23 51.22$\pm$0.32 82.47$\pm$0.02 81.89$\pm$0.04 73.55$\pm$0.02 73.78$\pm$0.07 84.77$\pm$0.00 85.18$\pm$0.01
V-A2 Whole-body Control
To validate the effectiveness of our proposed whole-body control training framework RoboDuet, we make a comparison with the Floating Base+IK. To be fair, we evaluate the solvability and workspace under the same error threshold by applying the average Euclidean distance error and quaternion geodesic distance error of RoboDuet as the pose error thresholds for IK solver. For both, the initial joint positions of the arm are set to the same as default positions mentioned in Section III-B1. According to Fig. 4, RoboDuet demonstrates a 12.18% improvement in solvability compared to the Floating Base+IK configuration. Furthermore, the workspace comparison suggests that whole-body control can significantly enhance the operational capabilities of the robotic arm. It is worth reiterating that solvability is defined as reaching the target pose from the current pose without any collisions throughout the entire process. Additionally, many samples are inherently unreachable due to the limitations imposed by the structure of the robotic arm itself.
TABLE IV: Workspace and Survival Rate of Combining the ARX5 with Different Legged Robots
@@ V-A2 Whole-body Control
Go1 0.8189 0.8518 96.60 99.91
Go2 0.8150 0.8019 96.56 91.69
A1 0.8305 0.7957 99.37 91.72
V-A3 Cross-Embodiment
To exhibit the cross-embodiment capability of RoboDuet, we select two additional quadruped robots Unitree A1 and Unitree Go2. We mount the ARX5 on these quadruped robots and train the system using the same method as in stage 1. After convergence, we obtain specific loco policies for the new embodiments. We evaluate the workspaces and survival rates under external forces by directly combining the previously trained arm policy for the Go1+ARX5 with newly developed locomotion policies. As shown in Table IV, even without additional training in stage 2, the newly integrated systems Go2+ARX and A1+ARX maintain relatively robust locomotion and accurate pose tracking, closely matching the performance of the original Go1+ARX system. This outcome underscores the exceptional zero-shot cross-embodiment capability of RoboDuet. Notably, under still conditions, A1+ARX achieves an even higher survival rate of 99.37% along with a larger workspace of 0.8305. Due to fundamental differences among the various embodiments, only Go1+ARX5 demonstrates an increase in stability and an expansion of workspace during motion through its custom-tailored cooperative policy. In contrast, the other two systems exhibit slight decreases in performance. Nevertheless, they still achieve a survival rate of nearly 91% under external perturbations.
@@ V-B Real-world Experiments
In terms of real-world experiments, we craft three distinct types of tasks to validate the effectiveness of our policy in the real world and its proficiency in handling diverse mobile manipulation challenges. We directly deploy Floating Base+IK and RobotDuet on a real-world robotic system for testing.
@@ V-B1 Extreme 6D Pose Tracking
To evaluate the robustness and generalization of end-effector pose tracking, we select five target poses to evaluate their self-collision and solvability. The target poses are illustrated in Fig. 5, with four representing extreme 6D target configurations. Poses 2 and 3 are close to the ground and intersect with the robot’s legs, while poses 4 and 5 are significantly elevated above the robot, beyond the training sampling range. Notably, pose 5 is located on the right side, nearing the robotic arm’s maximum reach, and oriented towards the robot’s head, representing an highly pathological configuration that is unsolvable. For each pose, we conduct eight repetitions, and the results are shown in Table V. The IK system fails to resolve poses 2, 4, and 5. However, RoboDuet maintains stability across all poses and approached optimal approximate solutions for physically unreachable poses. Although pose 3 can be solved using IK, the floating base policy lacks whole-body control, leading to self-collision issues. Additionally, as training did not include the situation that arm in motion, tracking during poses 4 and 5 results in a higher risk of tipping, which we regard as self-collision. Due to the self-collision penalty integrated into the training process and its ability to maintain locomotion while the arm is in motion, RobotDuet experiences no collisions in nearly 95% of trials, with the arm stopping effectively near the legs in poses 2 and 3.
@@ V-B2 Whole-body Control
To further analyze the impact of whole-body control on mobile manipulation, we design a set of tasks involving transfering a doll from varying heights, as shown in Fig. 6. For each height, we perform 10 repeated trials. A trial is considered successful only if the object is picked up and transferred to the next height without being dropped at any moment during this process. The system’s speed is controlled via a joystick on the remote controller, while the end-effector of the robotic arm is operated through a VR device.
TABLE VI: Success Rate of Floating Base + IK vs. RoboDuet
During the Transfer of Objects Between Different Heights
@@ 0 $\rightarrow$ 20 20 $\rightarrow$ 60 60 $\rightarrow$ 100 100 $\rightarrow$ 0
Floating 20% 40% 0% 0% 15.0%
Base+IK
RoboDuet 60% 80% 40% 50% 57.5%
The experimental results are shown in Table VI. When grasping objects on the ground, it is crucial to avoid both self-collisions and collisions with the ground. Lacking whole-body control capabilities, the Floating Base+IK approach is prone to self-collisions and failure, particularly when picking up objects positioned in front of or below the head. In contrast, RoboDuet can adjust its body posture to minimize collisions during grasping, resulting in a 40% increase in success rate. As the height increases, the need for body posture adjustments decreases, resulting in a 20% improvement in success rates for both approaches. Placing or retrieving the doll from the cup holder on a standing desk presented a particularly challenging task. The robotic arm was required to maintain a horizontal or top-down pose while operating near its maximum reach, when IK often fails to find a valid solution. RoboDuet demonstrats high solution stability and was able to reduce the vertical distance between the base and the standing desk by inclining its body, which successfully completed approximately 50% of the trials. In summary, across the complete object transfer process, RoboDuet achieved an average success rate of 57.5%, representing a 42.5% improvement over the Floating Base+IK.
V-B3 Cross-Embodiment
To evaluate the generalizability of our method across different robot embodiments, we directly deploy the policy trained on the Unitree Go1+ARX5 to the Unitree Go2+ARX5. The Unitree Go1 weighs 12 kg, while the Unitree Go2 weighs 18 kg, representing a 41.6% increase in mass. Despite the significant increase in mass, RoboDuet still demonstrated robust whole-body control across both platforms, as illustrated in Fig. 7. The figure depicts four key moments of the two systems tracking the same trajectory, demonstrating consistent 6D pose tracking and robust whole-body control capabilities across both configurations.

### v5

@@ IV-C Training Details
We train 4096 parallel agents using the IsaacGym simulator [20] and the Proximal Policy Optimization (PPO) algorithm [21]. We train all algorithms for 50,000 iterations across 3 seeds, with the two-stage training comprising 10,000 iterations for stage 1 and 40,000 for stage 2, employing the asymmetric actor-critic framework used in [22]. All neural networks are designed as the Multilayer Perceptron (MLPs) with ELU [23] activations for hidden layers. For the cooperative policy, the network architecture consists of three layers with 512, 256, and 128 neurons, respectively. In contrast to the cooperative policy, the unified policy merges feature extractors into one, so we double the number of neurons in the hidden layers. All training is performed on NVIDIA RTX 4090 GPUs.
@@ IV-E Metrics
To quantify the performance of the algorithms, we define several metrics: (1) velocity tracking error, measured by the difference between the target commands and actual states for $v_{x}$, $v_{y}$, and $\omega_{yaw}$; (2) position tracking error, determined by the error for $l$, $p$, $y$, and the Euclidean distance $D$; (3) orientation tracking error, calculated based on the angles $\alpha$, $\beta$, $\gamma$, and the quaternion geodesic distance $\zeta$; (4) survival rate, assessed by randomly applying forces of 10 to 20 newtons to the base for 2 seconds and calculating the proportion of robots that maintain a base height above 0.26 m across all samples; (5) solvability, defined as the ratio of collision-free poses to the total number of samples; and (6) workspace, defined as the area of the convex hull formed by all target commands within the tracking error threshold, ensuring no self-collisions throughout the tracking process. A sample is considered successful if the tracking of the end-effector pose meets the thresholds of $D\leq 0.03$ m for distance and $\theta\leq\pi/36$ for orientation, where $\theta$ is the angle between the two vectors obtained by applying the target orientation and end-effector orientation to the same unit vector. The maximum time allowed to reach the target command is 4 seconds, after which the average error over the following 2 seconds is computed.
@@ V-A Simulation Experiments
V-A1 Whole-body Control
To validate the effectiveness of our proposed whole-body control training framework RoboDuet, we make a comparison with the FB+IK. To be fair, we evaluate the solvability and workspace under the same error threshold by applying the average Euclidean distance error and the quaternion geodesic distance error of RoboDuet as the pose error thresholds for IK solver. For both, the initial joint positions of the arm are set to the same as default positions mentioned in Section III-B1. According to Fig. 4, RoboDuet demonstrates a 14.83% improvement in solvability compared to the FB+IK configuration. Furthermore, the workspace comparison suggests that whole-body control can significantly enhance the operational capabilities of the robotic arm. It is worth reiterating that solvability is defined as reaching the target pose from the current pose without any collisions throughout the entire process. Additionally, many samples are inherently unreachable due to the limitations imposed by the structure of the robotic arm itself.
Figure 4: Comparison of Solvability and Workspace between FB+IK and RoboDuet.
TABLE III: Ablation Results (scaled by $10^{-2}$) of Different Methods in Still and Move Modes
@@ V-A1 Whole-body Control
Metrics Still Move
Unified Two-Stage Cooperative RoboDuet Unified Two-Stage Cooperative RoboDuet
Velocity Tracking $\downarrow$ $v_{x}$ (m/s) 0.43$\pm$0.00 0.47$\pm$0.00 0.30$\pm$0.00 0.32$\pm$0.00 12.16$\pm$0.02 11.47$\pm$0.00 9.83$\pm$0.05 9.7$\pm$0.00
$v_{y}$ (m/s) 0.45$\pm$0.00 0.44$\pm$0.00 0.36$\pm$0.00 0.34$\pm$0.00 17.10$\pm$0.00 16.84$\pm$0.02 15.91$\pm$0.00 15.42$\pm$0.00
$\omega_{z}$ (rad/s) 0.37$\pm$0.00 0.35$\pm$0.00 0.33$\pm$0.00 0.32$\pm$0.00 61.83$\pm$0.03 62.06$\pm$0.03 60.82$\pm$0.05 60.59$\pm$0.02
Position Tracking $\downarrow$ $l$ (m) 4.25$\pm$0.04 4.13$\pm$0.00 1.90$\pm$0.02 1.97$\pm$0.00 4.01$\pm$0.01 4.25$\pm$0.01 1.88$\pm$0.00 1.91$\pm$0.02
$p$ (rad) 22.62$\pm$0.00 21.37$\pm$0.12 17.42$\pm$0.03 18.81$\pm$0.05 21.93$\pm$0.11 22.45$\pm$0.07 17.06$\pm$0.03 18.09$\pm$0.01
$y$ (rad) 12.01$\pm$0.01 11.52$\pm$0.04 7.94$\pm$0.00 7.43$\pm$0.00 12.15$\pm$0.00 11.43$\pm$0.00 7.70$\pm$0.00 7.22$\pm$0.00
$D$ (m) 14.47$\pm$0.01 13.84$\pm$0.01 10.79$\pm$0.01 11.08$\pm$0.00 13.56$\pm$0.02 13.05$\pm$0.01 10.58$\pm$0.00 10.75$\pm$0.00
Orientation Tracking $\downarrow$ $\alpha$ (rad) 44.15$\pm$0.03 47.76$\pm$0.03 40.46$\pm$0.05 39.17$\pm$0.02 42.50$\pm$0.08 47.63$\pm$0.07 43.22$\pm$0.04 39.89$\pm$0.01
$\beta$ (rad) 60.41$\pm$0.13 62.86$\pm$0.08 48.89$\pm$0.03 45.44$\pm$0.02 56.74$\pm$0.04 63.25$\pm$0.15 45.29$\pm$0.01 47.58$\pm$0.03
$\gamma$ (rad) 52.33$\pm$0.01 54.74$\pm$0.03 43.56$\pm$0.04 39.38$\pm$0.04 52.36$\pm$0.03 54.38$\pm$0.05 42.74$\pm$0.02 39.01$\pm$0.00
$\zeta$ (-) 51.15$\pm$0.01 53.46$\pm$0.02 48.18$\pm$0.00 47.14$\pm$0.00 50.58$\pm$0.00 53.03$\pm$0.01 48.36$\pm$0.00 47.53$\pm$0.00
Survival Rate (%) $\uparrow$ 87.35$\pm$0.78 95.66$\pm$0.06 94.49$\pm$0.03 98.20$\pm$0.01 93.03$\pm$1.34 99.09$\pm$0.01 98.56$\pm$0.02 99.96$\pm$0.00
Workspace ($m^{3}$) $\uparrow$ 51.36$\pm$0.46 60.10$\pm$0.05 84.08$\pm$0.06 85.79$\pm$0.02 78.96$\pm$0.04 81.30$\pm$0.02 86.71$\pm$0.01 87.05$\pm$0.01
V-A2 Ablation
To validate the significance of the two-stage training and the cooperative mechanism, which are the key components of RoboDuet, we conduct sufficient ablation experiments by comparing with Unified, Two-Stage and Cooperated. The results are shown in Table III, which demonstrate that all configurations meet the requirements for stable standing. Although two-stage training does not significantly improve end-effector pose tracking, it demonstrates a notable improvement in resisting external perturbations. Specifically, Two-Stage improves survival rates by 8.31% over the Unified in stationary standing, attributed to the focus of stage 1 on optimizing locomotion. On the other hand, Cooperated and RoboDuet outperform both the Unified and Two-Stage configurations across almost all metrics, indicating that the cooperative mechanism effectively decouples multi-task learning, enabling the two policies to collaborate while maintaining focus. Cooperated achieves the best performance in position tracking, largely due to its lack of the gait prior constraint from stage 1, allowing it to take a more aggressive approach in achieving higher tracking rewards which results a decrease on survival rates. In comparison, RoboDuet maintains comparable end-effector tracking performance while achieving robust locomotion motion ability and offers a larger operational workspace during motion. In summary, by effectively integrating cooperative policy and two-stage training, RoboDuet significantly enhances control performance in both tracking accuracy and gait stability, underscoring the essential role of these components.
TABLE IV: Workspaces and survival rates of zero-shot policy transfer from Go1 to Go2 and A1.
@@ V-A2 Ablation
Go1 0.8579 0.8705 98.20 99.96
Go2 0.8552 0.8216 96.32 93.21
A1 0.8416 0.8121 97.07 94.34
V-A3 Zero-shot Transfer
To assess the zero-shot transfer capability of RoboDuet, we introduce two additional quadruped robots, Unitree A1 and Unitree Go2, and mount the ARX5 on their backs. The policy trained on the Go1+ARX system is directly transferred to these newly integrated platforms. Subsequently, we evaluate their workspaces and survival rates in the presence of external disturbances, with the results presented in Table IV. Notably, under stationary standing conditions, Go2+ARX achieves a workspace of 0.8552, while A1+ARX attains a survival rate of 97.07%, closely aligning with the performance of the original Go1+ARX system. Due to fundamental differences among the various embodiments, only Go1+ARX5 demonstrates an increase in stability and workspace during motion through its custom-tailored cooperative policy. In contrast, the other two systems exhibit slight performance degradation. Nevertheless, they still achieve a survival rate of nearly 93% under external perturbations.
@@ V-B Real-world Experiments
In terms of real-world experiments, we design three distinct types of tasks to evaluate the effectiveness of our policy in the real world and its proficiency in handling diverse loco-manipulation challenges. We directly deploy FB+IK, AFB+ID, and RobotDuet on a real-world robotic system for testing.
@@ V-B1 Extreme 6D Pose Tracking
To evaluate the robustness and generalization of end-effector pose tracking in our method, we select five target poses to compare the self-collision and solvability between FB+IK and RoboDuet. The target poses are illustrated in Fig. 6, with four representing extreme 6D target configurations. Poses 2 and 3 are close to the ground and intersect with the robot’s legs, while poses 4 and 5 are significantly elevated above the robot, beyond the training sampling range. Notably, pose 5 is located on the right side, nearing the robotic arm’s maximum reach, and oriented towards the robot’s head, representing an highly pathological configuration that is unsolvable. For each pose, we conduct eight repetitions, and the results are shown in Table V. The IK system fails to resolve poses 2, 4, and 5. However, RoboDuet maintains stability across all poses and approached optimal approximate solutions for physically unreachable poses. Although pose 3 can be solved using IK, the floating base policy lacks whole-body control, leading to self-collision issues. Additionally, since the training does not include scenarios where the arm is in motion, tracking during poses 4 and 5 results in a higher risk of tipping, which we regard as self-collision. In contrast, due to the self-collision penalty integrated into the training process and its ability to maintain locomotion while the arm is in motion, RobotDuet experiences no collisions in nearly 95% of trials, with the arm stopping effectively near the legs in poses 2 and 3.
@@ V-B2 Whole-body Control
To further analyze the impact of whole-body control on loco-manipulation, we design a set of tasks involving transfering a doll from varying heights, as shown in Fig. 5. For each height, we calculate the average success rate from five repeated trials conducted by each of the three different operators. A trial is considered successful only if the object is picked up and transferred to the next height without being dropped at any moment during this process. The system’s speed is controlled via the right VR handle, while the end-effector is operated with the left handle.
TABLE VI: Success Rates Across Different Heights for Various Control Methods in Object Transfer Tasks.
@@ 0 $\rightarrow$ 20 20 $\rightarrow$ 60 60 $\rightarrow$ 100 100 $\rightarrow$ 0
FB+IK 3/15 7/15 0/15 0/15 10/60
AFB+ID 9/15 13/15 6/15 4/15 32/60
RoboDuet 10/15 12/15 8/15 9/15 39/60
The experimental results are shown in Table VI. When grasping objects on the ground, it is crucial to avoid self-collisions. Lacking whole-body control capabilities, the FB+IK approach is prone to self-collisions and failure, particularly when picking up objects positioned in front of or below the head. In contrast, AFB+ID and RoboDuet can adjust the body posture to minimize collisions during grasping, resulting in an almost 40% increase in success rate. However, since AFB+ID can only adjust the rotation around the roll axis based on interaction forces, it still struggles to grasp objects positioned directly in front of the head. As the height increases, the need for body posture adjustments decreases, resulting in a improvement in success rates for all approaches. Placing or retrieving the doll from the cup holder on a standing desk presented a particularly challenging task. The robotic arm was required to maintain a horizontal or top-down pose while operating near its maximum reach, when IK often fails to find a valid solution. AFB+ID relies on its gravity compensation mechanism and inverse dynamics solution, achieving a placement success rate of 40%. However, due to joint limit constraints, it can only successfully retrieve the doll when its ears or head are positioned near the outer edge of the cup holder. RoboDuet demonstrats high solution stability and is able to reduce the vertical distance between the base and the standing desk by inclining the body, which successfully completed approximately 55% of the trials. Compared to AFB+ID, RoboDuet achieves a 23% improvement in success rate for more challenging loco-manipulation tasks.
V-B3 Zero-shot Transfer
To evaluate the generalizability of our method across different robot embodiments, we directly deploy the policy trained on the Unitree Go1+ARX5 to the Unitree Go2+ARX5. The Unitree Go1 platform has a total mass of 18.35 kg, whereas the Unitree Go2 platform weighs 21.05 kg, representing a 14.7% increase. Despite the significant increase in mass, RoboDuet still demonstrated robust whole-body control across both platforms, as illustrated in Fig. 7. The figure depicts four key moments of the two systems tracking the same trajectory, demonstrating consistent 6D pose tracking and robust whole-body control capabilities across both configurations.